OpenStudy (calculusxy):

Point-slope form help! Write the point-slope form of the equation of the line through the given point with the given slope. through (-3,-2), slope = 1

OpenStudy (calculusxy):

I got y + 2 = x + 3.

OpenStudy (calculusxy):

(y - (-2)) = 1(x + 3) y + 2 = x + 3

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I would write it as `y + 2 = 1(x+3)` to show that the slope is m = 1. But you pretty much got it. Nice work.

OpenStudy (calculusxy):

thanks !

jimthompson5910 (jim_thompson5910):

you're welcome